OpenStudy (anonymous):

True or False: (1/2)(2n + 2) + -2 = n True False

OpenStudy (anonymous):

Let's do the math!\[\frac{1}{2}(2n+2)-2=n,\]\[n+1-2=n,\]\[n-1\neq n.\]So this is not true, as you can see.

OpenStudy (anonymous):

thank you.